Following are the annual returns on a portfolio over each of the last three years: Year...

  1. Following are the annual returns on a portfolio over each of the last three years:

Year

Return

2011

10%

2012

15%

2013

20%

  1. What was the arithmetic average return and geometric average return over the past three years?
  2. What would be your average return if you invested $200,000 per year at the beginning of each of the last three years?
  3. What would be your average return if you invested $100,000 at the beginning of 2011, $200,000 at the beginning of 2012 and $300,000 at the beginning of 2013?
Answer #1

A. Arthemathic Return= 15%
Geometric Return = 14.9275%

B. Average return will be 15%.
C. Average return will be 16.6662%

a. As themathic Returen - Simple Average retuen A Maw = 10+15+20 3 = 151 Geometric Retun - (3/(1+r) (1+r) (178) e) - 1 = ( 3ex Retuen 101. waghts 0.1866+ 1.667 10, 15 4.9995 0.3333 200,రా 10 6.5 201. 30, 16.66627. 60, on